OpenStudy (anonymous):

Determine the volume of the right circular cone generated by rotating the line y = 1/3x about the y-axis between y = 0 and y = 3. A) V = 81pi cu.units B) V = 84pi cu.units C) V = 82pi cu.units D) V = 85pi cu.units E) V = 83pi cu.units @Calculus1

OpenStudy (saifoo.khan):

A.

OpenStudy (saifoo.khan):

do u want the working?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok i'm assuming that it's y=x/3 not y = 1/(3x). Rotating around the y-axis; solve for x. y = x/3. x = 3y. Find x^2. x^2 = 9y^2. Integrating it using the limits of 0 and 3, remember to add the pi and integrate 9y^2. It comes out to be: 81pi cubic units.

OpenStudy (anonymous):

A is correct.

OpenStudy (anonymous):

Thanks a lot

OpenStudy (saifoo.khan):